OpenStudy (anonymous):

Hey @satellite73 I have another kind of problem for you. If possiable. 7x^2/√x^5

OpenStudy (anonymous):

\[\frac{7x^2}{\sqrt{x^5}}\]?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

ok denominator is \[\sqrt{x^5}=x^2\sqrt{x}\]

OpenStudy (anonymous):

k so you just used the thing you taught me last night right.

OpenStudy (anonymous):

now you can rewrite it as \[\frac{7x^2}{x^2\sqrt{x}}\] yup, but we are not done yet

OpenStudy (anonymous):

the 2 goes into 5, two times with the remainder of 1.

OpenStudy (anonymous):

yes, exactly

OpenStudy (anonymous):

k

OpenStudy (anonymous):

that is how we go from \[\sqrt{x^5}\] to \[x^2\sqrt{x}\]

OpenStudy (anonymous):

k.

OpenStudy (anonymous):

but there is more to do you have \[\frac{7x^2}{x^2\sqrt{x}}\]

OpenStudy (anonymous):

you can cancel an \(x^2\) top and bottom to get \[\frac{7x^2}{x^2\sqrt{x}}=\frac{7}{\sqrt{x}}\]

OpenStudy (anonymous):

and finally, if you want to write this in "simplest radical form" i.e. with no radical in the denominator, multiply top and bottom by \(\sqrt{x}\) to get \[\frac{7}{\sqrt{x}}=\frac{7}{\sqrt{x}}\times \frac{\sqrt{x}}{\sqrt{x}}=\frac{7\sqrt{x}}{x}\]

OpenStudy (anonymous):

some people don't like radicals in the denominator, some don't care. either way it is the same

OpenStudy (anonymous):

hmmm.

OpenStudy (anonymous):

up to your teacher really

OpenStudy (anonymous):

for example \[\frac{1}{\sqrt{2}}=\frac{1}{\sqrt{2}}\times \frac{\sqrt{2}}{\sqrt{2}}=\frac{\sqrt{2}}{2}\] just two different forms of the same number

OpenStudy (anonymous):

Well the answers have radicals in them. sooo